An open flask contains air at 27°C. To what temperature at which 2/5 th mass of air escapes the flask
1
Expert's answer
2019-04-26T11:34:06-0400

According to Charle's law,


"\frac {V_1} {T_1}=\frac {V_2} {T_2} (1)"

Using (1) we got:


"T_2=\frac {T_1 \times V_2} {V_1} (2)"

In our case,

T1=300 K,


"V_1=\frac {3V_2} {5}"

We got:

T2=500 K or 227 °C


Answer:

500 K or 227 °C